drift velocities

From: Ian Bearden (bearden@nbi.dk)
Date: Tue Jan 14 2003 - 09:13:22 EST

  • Next message: Eun-Joo Kim: "Re: drift velocities"
    Can someone please remind me where the drift velocities are stored?
    I would really like to know what they are, but cannot seem to find the
    information.
    Please help!
    Cheers,
    Ian
    
     --------------------------------------------------------------------
    | I.G. Bearden                                                       |
    | Niels Bohr Institute Tb 3   email:      bearden@nbi.dk             |
    | Blegdamsvej 17              phone:      (+45) 35 32 53 23          |
    | København Ø                   FAX:      (+45) 31 42 10 16          |
    | Danmark                       Www:      http://www.nbi.dk/~bearden |
     --------------------------------------------------------------------
    
    -----Original Message-----
    From: owner-brahms-l@bnl.gov [mailto:owner-brahms-l@bnl.gov]On Behalf Of
    Apache
    Sent: 14. januar 2003 14:42
    To: brahms-l@bnl.gov; olchansk@bnl.gov; videbaek@bnl.gov
    Subject: Shift report 20030114 00:00-08:00
    
    
    
    ------------------------------------------------------------
    Date: Tue Jan 14 08:41:40 2003, Shift: 00:00-08:00
    Supervisor: JIJ, with: BB
    Log book#13 pages 59 - 59
    
    Shift Summary:
    ------------------------------------------------------------
    1.
    Long store, usable beam from midnight until 6:30, low background, i.e. H1
    and H2 scalers with rates that decreased steadily from 20k/s to 3-4k/s.
    
    2.
    7367 - 56k trigger 6, D1 @ 1/4 A-pol, others off
    7368 - 59k trigger 6, FS @ 1/4 B-pol, D5 off
    7369 - 47k trigger 6, same setting
    7370 -  6k trigger 6, same setting
    7371 - pedestal run
    7372 - TPC pedestal run
    7373 - DVM run
    
    In the course of run 7366-7370 a total of 17.7k trigger 3 events were
    recorded with D5 off. Another 2.3k trigger 3 needed for goal at 20k.
    
    3.
    The 4 suspicious top channels of H2 mentioned previously have currently
    demand voltage of -1400V. This gives read-back varying between -1575V
    and -1625V, although we could not see any periodicity in these variations.
    Tried to look at differences in ADC gains for previous runs, but
    interpretation is hard as also the good channels vary in similar manner as
    the bad ones. Some more testing should be performed on the power supply
    during access Wednesday.
    
    Horizontal fibers in FS were turned on this morning at 08:15. They have been
    off until now, but were sadly missed in Copenhagen.
    
    TPM1 still with anodes at 1160V until further notice. T3 still off.
    
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Jan 14 2003 - 09:15:12 EST